ADNAN MUHAMMED

Grade 12,

A car initially at rest starts moving on a straight line with constant acceleration and attains speed of 16m/s in time t s.Average speed of the car in this time interval is

A car initially at rest starts moving on a straight line with constant acceleration and attains speed of 16m/s in time t s.Average speed of the car in this time interval is

Grade:11

1 Answers

Vikas TU
7 years ago
From first law of equation,
v = u + at
16 = 0 + at
a = 16/t m/s^2
 
Distance covered in t seconds,
s = 0 + 0.5*16/t * t^2 => 8t meter.
Average speed = tota distance / total time => 8t/t = > 8 m/s.